the sum of the fourth and twelfth term of an A.P is 120.what is the sum of the first 15 terms of the A.p

Answer by richwmiller(17219)

an = a + (n - 1)d
a4 = a + 3d
a12 = a + 11d
120=2a+14d
S = ½(2a + (n-1)d)n
S = ½(2a + (14)d)15
S = ½(120)15
S=900